The afternoon shift was devoted to a differential tuning of the DASes. Details on that in a separate entry, here we report about the ISC/Automation part:
- we relocked with no particular difficulties, also the DRMI locked rather easily;
 - again we had to tune a bit the PRCL gains along the way, but to a much lesser extent with respect to yesterday;
 - we unlocked once at 17:09:17 at STEP3, the unlocks seemed to originate from INJ (Figure 1);
 - we then locked at CARM_NULL_1F at 17:20:34 UTC; we noticed that PSTAB was still closed, causing the low-frequency oscillations on the IMC_TRA and related powers; PSTAB was then opened and left open, to the benefit on the stability of the lock;
 - fixing the WI timing cured the DIFFp issues;
 - then, while waiting for the thermalization needed for the TCS activity, we tuned more properly the demodulation phases for the SR DDemod signals; signals were behaving correctly already, but we improved them a bit; SR_TX gain had been reverted to 0.5 as previously written, but TY is still 0.2; to be checked if it can be raised back to 0.5;
 - SRCL_SET (with target OS to 4 Hz) was around 15 nm;
 - at around 19:10 UTC the TCS step was done, and the lock kept until 20:55:05 UTC, when we unlocked by hand in order have a relock from scratch in the new TCS conditions;
 - we relocked up to the DRMI and we immediately unlocked by hand, in order to test a new feature of the automation for the absorption measurement: now, in addition to keeping aligned PR and SR, ITF_LOCK asks the arms to go to the Single Bounce configuration, the proper one being determined by which HWS is targeted to be used from the TCS_MAIN.ini configuration file;
 - we did then a full relock to CARM_NULL_1F, reached at 21:21:24 UTC;
 - after 45 minutes we set up the absorption measurement, with the following parameters:	
- HWS: DET;
 - 60 minutes to reach the full thermalization of the ITF;
 - HWS initialization and 20 minutes of reference data while locked;
 - unlock with CH kept off, PR and SR kept aligned, and Single Bounce from NI (done with ARMS_LOCK directly); 7200s of measurement;
 - CH turned back on, another 7200s;
 - when ABSORPTION_MEAS_DONE will be reached and the timeout passed, the relock can happen; the keep_aligned flag should be put back to False, but it will not prevent the lock acquisition (only notifications will appear, and the PR/SR/ARMS alignment will be performed before the next lock acquisition attempt, instead of immediately after the unlock.
 
 
We leave the ITF locked, with the measurement in progress.
The ITF relock can be performed not before approximately 05:00 LT. The keep_aligned flag has been already reverted to False.